How close to the right edge of the 56 kg picnic ta
Answer #1

moment of table = moment of person
56kg * (2.10m/2 - 0.55m) * g = 90kg * (0.55m - x) * g
56 * 0.5 = 49.5 - 90x
x = 0.23 m =0.019 cm